Give Big Θ (tight Big-Oh) expressions for each problem. For full credit, they must be fully simplified. k-1 2 k-1
Answer #1

`Hey,

Note: Brother in case of any queries, just comment in box I would be very happy to assist all your queries

1) sum(k/2)=n*(n+1)=O(n^2)

2) We can approximate inner summation as sum(k^3). So, final would be of order O(n^4)

3) 2+2^2+2^3....2^n

S=2*(2^n-1)

So, S=O(2^n)

Kindly revert for any queries

Thanks.